A total of $4300 was invested, part of it at 5% interest and the remainder at 10%. If the total yearly interest amounted to $310, how much was invested at each rate?
Answer by mananth(16949)

total of $4300 was invested,
Let $x be invested at 5% interest
remainder $4300-x at 10%.
total yearly interest = $310
..
0.05x+0.1(4300-x)=310
0.05x+430-0.1x=310
multiply the equation by 100
5x+43000-10x=31000
-5x=-12000
x= 2400
$ 2400 invested at 5%
Balance (4300-2400) = $1900 at 10%